### v1

@@ 3 An Efficient Newton-type Multilevel Method
The ML-Newton method presented in the previous section chooses the coarse level for constructing the associated coarse direction according to (10). Although, convergence to the unique minimizer can now be established due to (10), the local rates for multilevel methods, whether Newton-based or otherwise, are generally linear-quadratic. Indeed, without additional assumptions, linear-quadratic rates are the best one can typically expect.
@@
Lemma 3.1.
Let $x_{k+1}=x_{k}+d_{m,k}$, where $d_{m,k}$ as in (13). If inequality (14) holds, then $\|\nabla f(x_{k})\|=0\iff\|R_{i}\nabla f(x_{k})\|=0$ for all $k\in\mathbb{N}$ and $1\leq i\leq m-1$.
Proof 3.2.
Let us fix $k\in\mathbb{N}$ and $1\leq i\leq m-1$. The implication $\Longrightarrow$” holds immediately. Assume now that $\|R_{i}\nabla f(x_{k})\|=0$. Then $R_{i}\nabla f(x_{k})=0$ and $x_{k+1}=x_{k}$. Hence, $\|R_{i}\nabla f(x_{k})\|=\|R_{i}\nabla f(x_{k+1})\|=0$ and $\|\nabla f(x_{k})\|=\|\nabla f(x_{k+1})\|$. We will prove the result by contraposition. If $\|\nabla f(x_{k})\|\neq 0$, then $\|\nabla f(x_{k+1})\|\neq 0$. However, putting this all together, the condition (14) is violated since $\sigma\in(0,1]$. Hence, $\|\nabla f(x_{k})\|=0$ as required.
Lemma 3.1 shows that $d_{m,k}$ is not only effective at iteration $k+1$, as it may initially appear, but also at iteration $k$. Notice also that $\|R_{i}\nabla f(x_{k})\|>\varepsilon$ is no longer required in our setting since the multilevel method can now reach very high accuracy. Consequently, the condition (14) can effectively replace the classical condition (10), while still ensuring progress of the multilevel scheme at each iteration.
@@
(15) $x_{k+1}=x_{k}+d_{m,k}.$
If there exists no $1\leq i\leq m-1$ such that (14) is satisfied, then we set $i=m$ and $R_{m}=I_{n}$ and perform pure Newton steps, i.e.,
@@
0: initial value $x_{0}\in\mathbb{R}^{n}$, restriction operators $R_{i,k}$, $\sigma\in(0,1]$
@@
2:  Compute $i^{*}=\min\left\{1\leq i\leq m:\|R_{i,k}\nabla f(x_{k+1})\|\geq\sigma\|\nabla f(x_{k+1})\|\right\}$
@@
4:   $x_{k+1}=x_{k}-P_{i^{*},k}\left(R_{i^{*},k}\nabla^{2}f(x_{k})P_{i^{*},k}\right)^{-1}R_{i^{*},k}\nabla f(x_{k})$
@@
Furthermore, one could define iteration-dependent restriction and prolongation operators, i.e., $R_{i,k}$ and $P_{i,k}$, respectively, and derive the algorithm’s coarse steps without modifying the above analysis. A generic Adaptive Multilevel Newton method (AML–Newton) with iteration-dependent operators is presented in Algorithm 1. A process for computing $i^{*}$ in step 2 in Algorithm 1 will be discussed in Section 5.1. In any case though, one has to compute a trial step to evaluate (14). This may come with the cost of constructing extra coarse directions at each iteration with cost $\mathcal{O}\big(n_{i}^{3}\big)$, and thus, per iteration, Algorithm 1 is more expensive than the ML–Newton method. The total complexity of Algorithm 1 is difficult to measure since it is very much dependent on the design of coarse models and the choice of $\sigma$. In the worst scenario though the method will compute m1m-1 coarse steps plus a Newton step. In Section 5.1 we will discuss how to design coarse models such that in practice at most two coarse steps with $n_{i}\ll n$ are computed on average, at each iteration.
We also note that the idea of taking a trial step is not new in optimization. It has been a key component in trust-region methods [7]. Nevertheless, although the main idea looks the same, it is in fact different from that of Algorithm 1. In trust region methods, one looks at the ratio of the actual vs. predicted decrease of the objective function to decide whether a candidate step is accepted. Then, the trust-region radius is adjusted accordingly. Here, instead, we look at the ratio of gradient norms to, in essence, determine whether a level is accepted. As we will see later, the norm of the reduced gradient decreases faster than that of the true gradient, thus our goal is to find a level that imposes a similar decrease to the latter norm.
@@ 4 Convergence Analysis
The global convergence behavior of multilevel Newton methods for strictly convex functions has been thoroughly analyzed and shown to align with that of the classical Newton method (see, e.g., [17, 36, 33]). However, its local convergence analysis remains weaker compared to the well-established local behavior of the classical Newton method. In this section, we fill this gap by presenting a local convergence analysis of Algorithm 1, establishing its quadratic rate under two settings:
@@
    self-concordant functions.
@@
where $x_{k}$ arises from Algorithm 1.
@@ 4.1 Strongly Convex Functions
In this section we analyze Algorithm 1 for strongly convex functions with Lipschitz Hessians [3, 27]. That is, we assume there exist constants $\mu,L>0$ such that
@@
We note that the quadratic convergence rate for ML–Newton in the subspace (see (29)) is not new. Similar results appear in [17]. Our proof, however, is different and more general, as it supports iteration-dependent operators.
@@ 4.2 Self-concordant Functions
Self-concordance is crucial when analyzing Newton’s method as it offers local convergence rates that are invariant to the affine transformation of variables. These rates are important, and are on par with the Newton’s method’s iterations; it seems that the class of self-concordant functions captures effectively the convergence behavior of Newton’s method observed in practical applications. In this section, in a similar fashion, we attempt to explain our algorithm’s practical behavior using self-concordant functions. In particular, we will show a local quadratic convergence rate that it is independent of unknown constants such as $\mu,L$ involved in Theorem 4.3.
@@
The above quantity is well-defined due to $R_{i,k}\nabla^{2}f(x)P_{i,k}\succ 0$. When setting $x\coloneq x_{k}$, it coincides with the approximate decrement $\hat{\lambda}_{i,k}$ used for the analysis of past multilevel algorithms [17, 33, 36]. It represents the local norm of the reduced gradient induced by the reduced Hessian matrix at iteration $k$ and level $i$. In what follows we collect results for self-concordant functions that will be useful for establishing the convergence results. Let $r\coloneq\|y-x\|_{x}$, where $\|\cdot\|_{x}$ is defined in (4). Given (3), for any $x,y\in\mathbb{R}^{n}$ such that $Mr<1$, it is possible to show [27] that
@@
The following result, proved in [33, Lemma 2.3], shows that the approximate decrement is always smaller than the Newton decrement for all $k\in\mathbb{N}$, i.e.,
@@
For our analysis we will need also the following lemma, proved in [34].
Lemma 4.7 (Lemma 3.1 in [34]).
@@ 4.3 Discussion of the Convergence Results
We will finish this section by comparing Algorithm 1 and Newton’s method. Note that the only difference between the estimates (52) and Newton’s estimates in (2) and (6) is the presence of $0<\sigma\leq 1$ in (52). Thus, Newton’s local quadratic rate is faster than that of Algorithm 1. Since, however, quadratic rates are very fast, the main difference between Algorithm 1 and Newton’s method comes from the region of their local convergence. The estimates (52) imply a restricted region for the local quadratic convergence, and thus Algorithm 1 requires more steps in the first phase than Newton’s method. However, this is not necessarily a disadvantage of Algorithm 1 because it generates fast iterates in levels with $n_{i}<n$ or even $n_{i}\ll n$. On the other hand, it is known that Newton’s global convergence achieves a linear rate [21], and thus it does not justify its high computational demands. Therefore, one should find a replacement of Newton’s method in the first phase and only apply Newton in the second phase. This is exactly what Algorithm 1 does, and importantly, through condition (30), it offers an automatic way to switch from linear to a Newton-like quadratic convergence in the late stages of optimization. We will verify the theoretical findings of this section in Section 5.
@@ 4.4 Relation to Randomized Subspace Newton Method
Newton-type subspace methods have attracted a lot of interest due their efficiency for large-scale optimization [11, 16, 35]. By keeping the subspace fixed, the key difference to multilevel methods is that they can converge to the minimizer of $f$ without ever computing directions on the fine level. This fact makes subspace methods superior to multilevel methods in practical applications where $n$ is prohibitively large to form the Hessian matrix. Regarding their convergence rate, they achieve probabilistic linear convergence rates, unless extra assumptions related to the problem structure are imposed.
@@
Lemma 4.14 ([19, 20]).
@@
holds with probability at least $(1-\delta)^{2}$ for each $k$. That is, at each iteration, the subspace method advances to $k+1$ quadratically with high probability. More importantly, it implies high total probability of converging quadratically. In particular, suppose that the algorithm takes $K_{\varepsilon}$ steps in the the second phase. Then, the total probability to converge quadratically is $(1-\delta)^{2K_{\varepsilon}}$. Since quadratic convergence is very fast ($K_{\varepsilon}$ is at most $5$ or $6$ in practice) and $\delta$ is small by Theorem 4.15, then the total probability $(1-\delta)^{2K_{\varepsilon}}$ remains high. Particular examples of $\mathcal{D}$ that satisfy Lemma 4.14 include but are not limited to Gaussian distributions, e.g., Gaussian matrices, and uniform distributions, e.g., 1- and s-hashing matrices, for more details see [5] and references therein. Explicit convergence rates with these particular examples can be derived in a similar manner, nevertheless, the theorem does not seem practical since quadratic rates in the subspace could not be observed (which is a requirement in our theory, see Remark 4.5). However, Theorem 4.15 is important as it demonstrates that probabilistic quadratic rates are possible. Specifically, it suggests that, for such rates to be attainable, one should design restriction operators $R_{k}$ that satisfy Lemma 4.14 and ensure that the quadratic rate is activated within the relevant subspace. A probabilistic analysis for the class self-concordant functions is not straightforward as it requires a similar result to Lemma 4.14 based on local norms.
@@ 5 Numerical Experiments
    that Algorithm 1 can attain quadratic convergence with a total complexity lower than that of the Subspace Newton method [11].
@@
Both problems are widely used in practical machine learning applications. The logistic loss satisfies the strongly convex assumption under mild conditions [31], whereas the Poisson loss is a self-concordant function. Table 1 shows details of the datasets used and the associated loss function.
@@ 5.1 Implementation Details
Note that, under Definition 5.1, we have $\omega=\xi=1$ and thus the convergence rate for strongly convex functions is simplified as in (52). From now onward, unless otherwise specified, when Algorithm 1 uses fixed operators, the former definition of the $S_{n_{i}}$ subsets will be used to construct $R_{i}$ in Definition 5.1. We call this version of Algorithm 1 AML–Newton (fixed) as it uses fixed subspaces. If Algorithm 1 performs steps using iteration-dependent $R_{i,k}$, then they are generated based on the latter definition of the $S_{n_{i}}$ subsets. We call this version of Algorithm 1 AML–Newton (random) since the subspaces are generated randomly at each iteration.
It is well known that multilevel (or subspace) methods are particularly effective for problems that have a low-rank structure, e.g., their Hessian is nearly low-rank, or the second-order information is concentrated in the leading eigenvalues [17, 33]. In contrast, when the curvature is spread across many directions, the first phase of multilevel methods can be slow, which may prevent the algorithm from entering the quadratic phase. In this case, larger values of $\sigma$ should be selected in order allow for more Newton steps. Similarly, one could randomly permute the $m$ levels at each iteration, introducing an approximately $1/m$ probability of selecting the Newton step. If the problem exhibits a low-rank structure, we wish the coarse models to be employed across the first phase, and hence $\sigma$ should be selected small, nevertheless not too small to compromise the quadratic second stage of the algorithm.

### v2

@@ 3 An Efficient Newton-type Multilevel Method
The multilevel Newton method presented in the previous section chooses the coarse level for constructing the associated coarse direction according to (10). Although, convergence to the unique minimizer can now be established due to (10), the local rates for multilevel methods, whether Newton-based or otherwise, are generally linear-quadratic. Indeed, without additional assumptions, linear-quadratic rates are the best one can typically expect.
@@
(15) $x_{k+1}=x_{k}+d_{m,k},$
otherwise we check condition Eq. 14 for $i+1$. If there exists no $1\leq i\leq m-1$ such that (14) is satisfied, then we set $i=m$ and $R_{m}=I_{n}$ and perform pure Newton steps, i.e.,
@@
0: initial value $x_{0}\in\mathbb{R}^{n}$, restriction operators $R_{i,k}$, damping parameters $\alpha_{k},\sigma\in(0,1]$
@@
2:  Compute
$i^{*}:=\min\left\{\,1\leq i\leq m:\;\begin{aligned} \|R_{i,k}\nabla f(x_{k})\|&\geq\sigma\|\nabla f(x_{k})\|,\\ \|R_{i,k}\nabla f(x_{k+1})\|&\geq\sigma\|\nabla f(x_{k+1})\|\end{aligned}\right\}.$
@@
4:   Set $x_{k+1}:=x_{k}-\alpha_{k}P_{i^{*},k}\left(R_{i^{*},k}\nabla^{2}f(x_{k})P_{i^{*},k}\right)^{-1}R_{i^{*},k}\nabla f(x_{k})$.
@@
Furthermore, one could define iteration-dependent restriction and prolongation operators, i.e., $R_{i,k}$ and $P_{i,k}$, respectively, and derive the algorithm’s coarse steps without modifying the above analysis. A generic Adaptive Multilevel Newton method (AML–Newton) with iteration-dependent operators and damping parameter $\alpha_{k}$ is presented in Algorithm 1. Note that, when applying iteration-dependent operators, Algorithm 1 must also check the condition $\|R_{i,k}\nabla f(x_{k})\|\geq\sigma\|\nabla f(x_{k})\|$ at each iteration. We show below that, in the classical multilevel setting with fixed operators, this condition follows directly from Eq. 14 and can therefore be removed from step 2. There are several ways to compute the damping parameter at each iteration. The Armijo rule is often preferred over other strategies in second-order methods because it is robust and automatically adapts the step size, accepting the full Newton step when it is safe.
A process for computing $i^{*}$ in step 2 in Algorithm 1 will be discussed in Section 5.1. In any case though, one has to compute a trial step to evaluate (14). This may come with the cost of constructing extra coarse directions at each iteration with cost $\mathcal{O}\big(n_{i}^{3}\big)$, and thus, per iteration, Algorithm 1 is more expensive than the ML–Newton method. The total complexity of Algorithm 1 is difficult to measure since it is very much dependent on the design of coarse models and the choice of $\sigma$. In the worst scenario though the method will compute m1m-1 coarse steps plus a Newton step. In Section 5.1 we will discuss how to design coarse models such that in practice at most two coarse steps with $n_{i}\ll n$ are computed on average, at each iteration.
We also note that the idea of taking a trial step is not new in optimization. It has been a key component in trust-region methods [conn2000trust]. Nevertheless, although the main idea looks the same, it is in fact different from that of Algorithm 1. In trust-region methods, one looks at the ratio of the actual vs. predicted decrease of the objective function to decide whether a candidate step is accepted. Then, the trust-region radius is adjusted accordingly. Here, instead, we look at the ratio of gradient norms to, in essence, determine whether a level is accepted. As we will see later, the norm of the reduced gradient decreases faster than that of the true gradient, thus our goal is to find a level that imposes a similar decrease to the latter norm.
@@ 4 Convergence Analysis
We start by discussing the global convergence of Algorithm 1. To show global convergence we consider two cases: (i) the restriction operator is fixed at each iteration, i.e, $R_{i,k}\equiv R_{i}$, (ii) it varies as in Algorithm 1.
Let $R_{i}$, $1\leq i\leq m-1$, be fixed and assume for simplicity that the coarse step is always accepted (if no coarse model accepted at an iteration $k$, then the method progresses as Newton’s method whose global convergence is already established). Then by inequality Eq. 14 there exists a $j\in\{1,\ldots,m-1\}$ with
$\|R_{j}\nabla f(x_{k})\|\geq\sigma\|\nabla f(x_{k})\|.$
Thus, the method takes effective steps at each iteration similar to classical Newton multilevel method (see condition (10)). Then, assuming that the function is strongly convex with Lipschitz continuous gradients and Hessians the required global result follows from [ho2019newton]. It can be shown that
(17) $f(x_{k+1})\leq f(x_{k})-\nu_{k},$
where $\nu_{k}\coloneq\sigma C\|\nabla f(x_{k})\|^{2}$ for some $C>0$ and for all $k\in\mathbb{N}$ and $x_{0}\in\mathbb{R}^{n}$. Specifically, the authors show that (17) remains true even when the damping parameter $\alpha_{k}$ arises from the Armijo rule [ho2019newton, Lemma 3.1] and that the Armijo rule accepts the unit step [ho2019newton, Lemma 3.8]. Analogously, we can show similar global results for Algorithm 1 when assuming strictly convex self-concordant functions using Lemmas 3.1 and 3.4 proved in [tsipinakis2024multilevel]. In this case $\nu_{k}$ depends on the Newton decrement instead of norm of the gradient (see [tsipinakis2024multilevel]).
If Algorithm 1 is applied with iteration-dependent operators, then Eq. 14 does not necessarily imply $\|R_{i,k}\nabla f(x_{k})\|\geq\sigma\|\nabla f(x_{k})\|$ anymore, as $\|R_{i,k}\nabla f(x_{k})\|$ can be arbitrarily small compared to $\|\nabla f(x_{k})\|$. Thus, in order to establish the global convergence of the algorithm we need to check $\|R_{i,k}\nabla f(x_{k})\|\geq\sigma\|\nabla f(x_{k})\|$ at each iteration alongside Eq. 14. Then, as previously, for both strongly convex and self-concordant functions an inequality such as Eq. 17 can be shown from [ho2019newton] and [tsipinakis2024multilevel].
The above discussion shows that the global convergence behavior of the multilevel Newton methods for strictly convex functions aligns with that of the classical Newton method, (similar reductions in the value of the objective function have been shown in [nesterov2018lectures, boyd2004convex]). However, their local convergence analyses remain weaker compared to the well-established local behavior of the classical Newton method. In this section, we fill this gap by presenting a local convergence analysis of Algorithm 1, establishing its quadratic rate under two settings:
@@
    strictly convex self-concordant functions.
@@
where $x_{k}$ arises from Algorithm 1. For the remainder of this section, to derive the local convergence rates of Algorithm 1, we set $\alpha_{k}\equiv 1$ By the discussion above, if the current iterate is sufficiently close to the minimizer, the Armijo rule will always accept the unit step.
@@ 4.1 Strongly Convex Functions
In this section we analyze Algorithm 1 for strongly convex functions with Lipschitz Hessians [boyd2004convex, nesterov2018lectures]. That is, we assume there exist constants $\mu,L>0$ such that
@@
We note that the quadratic convergence rate for ML–Newton in the subspace (see (30)) is not new. Similar results appear in [ho2019newton]. Our proof, however, is different and more general, as it supports iteration-dependent operators.
@@ 4.2 Self-concordant Functions
Self-concordance is crucial when analyzing Newton’s method as it offers local convergence rates that are invariant to the affine transformation of variables. These rates are important, and are on par with the Newton’s method’s iterations; it seems that the class of self-concordant functions captures effectively the convergence behavior of Newton’s method observed in practical applications. In this section, in a similar fashion, we attempt to explain our algorithm’s practical behavior using strictly convex self-concordant functions. In particular, we will show a local quadratic convergence rate that it is independent of unknown constants such as $\mu,L$ involved in Theorem 4.3.
@@
The above quantity is well-defined due to $R_{i,k}\nabla^{2}f(x)P_{i,k}\succ 0$. When setting $x\coloneq x_{k}$, it coincides with the approximate decrement $\hat{\lambda}_{i,k}$ used for the analysis of past multilevel algorithms [ho2019newton, tsipinakis2024multilevel, tsipinakis2023lowrank]. It represents the local norm of the reduced gradient induced by the reduced Hessian matrix at iteration $k$ and level $i$. In what follows we collect results for self-concordant functions that will be useful for establishing the convergence results. Let $r\coloneq\|y-x\|_{x}$, where $\|\cdot\|_{x}$ is defined in (4). Given (3), for any $x,y\in\mathbb{R}^{n}$ such that $Mr<1$, it is possible to show [nesterov2018lectures] that
@@
The following result, proved in [tsipinakis2024multilevel, Lemma 2.3], shows that the approximate decrement is always smaller than the Newton decrement for all $k\in\mathbb{N}$, i.e.,
@@
For our analysis we will need also the following lemma, proved in [tsipinakis2025convergence].
Lemma 4.7 (Lemma 3.1 in [tsipinakis2025convergence]).
@@ 4.3 Discussion of the Convergence Results
We will finish this section by comparing Algorithm 1 and Newton’s method. Note that the only difference between the estimates (53) and Newton’s estimates in (2) and (6) is the presence of $0<\sigma\leq 1$ in (53). Thus, Newton’s local quadratic rate is faster than that of Algorithm 1. Since, however, quadratic rates are very fast, the main difference between Algorithm 1 and Newton’s method comes from the region of their local convergence. The estimates (53) imply a restricted region for the local quadratic convergence, and thus Algorithm 1 requires more steps in the first phase than Newton’s method. However, this is not necessarily a disadvantage of Algorithm 1 because it generates fast iterates in levels with $n_{i}<n$ or even $n_{i}\ll n$. On the other hand, it is known that Newton’s global convergence achieves a linear rate [karimireddy2018global], and thus it does not justify its high computational demands. Therefore, one should find a replacement of Newton’s method in the first phase and only apply Newton in the second phase. This is exactly what Algorithm 1 does, and importantly, through condition (31), it offers an automatic way to switch from linear to a Newton-like quadratic convergence in the late stages of optimization. We will verify the theoretical findings of this section in Section 5.
@@ 4.4 Relation to Randomized Subspace Newton Method
Newton-type subspace methods have attracted a lot of interest due their efficiency for large-scale optimization [gower2019rsn, hanzely2020stochastic, tsipinakis2025regularized]. By keeping the subspace fixed, the key difference to multilevel methods is that they can converge to the minimizer of $f$ without ever computing directions on the fine level. This fact makes subspace methods superior to multilevel methods in practical applications where $n$ is prohibitively large to form the Hessian matrix. Regarding their convergence rate, they achieve probabilistic linear convergence rates, unless extra assumptions related to the problem structure are imposed.
@@
Lemma 4.14 ([johnson1984extensions, kane2014sparser]).
@@
holds with probability at least $(1-\delta)^{2}$ for each $k$. That is, at each iteration, the subspace method advances to $k+1$ quadratically with high probability. More importantly, it implies high total probability of converging quadratically. In particular, suppose that the algorithm takes $K_{\varepsilon}$ steps in the the second phase. Then, the total probability to converge quadratically is $(1-\delta)^{2K_{\varepsilon}}$. Since quadratic convergence is very fast ($K_{\varepsilon}$ is at most $5$ or $6$ in practice) and $\delta$ is small by Theorem 4.15, then the total probability $(1-\delta)^{2K_{\varepsilon}}$ remains high. Particular examples of $\mathcal{D}$ that satisfy Lemma 4.14 include but are not limited to Gaussian distributions, e.g., Gaussian matrices, and uniform distributions, e.g., 1- and s-hashing matrices, for more details see [cartis2022randomised] and references therein. Explicit convergence rates with these particular examples can be derived in a similar manner, nevertheless, the theorem does not seem practical since quadratic rates in the subspace could not be observed (which is a requirement in our theory, see Remark 4.5). However, Theorem 4.15 is important as it demonstrates that probabilistic quadratic rates are possible. Specifically, it suggests that, for such rates to be attainable, one should design restriction operators $R_{k}$ that satisfy Lemma 4.14 and ensure that the quadratic rate is activated within the relevant subspace. A probabilistic analysis for the class self-concordant functions is not straightforward as it requires a similar result to Lemma 4.14 based on local norms.
@@ 5 Numerical Experiments
    that Algorithm 1 can attain quadratic convergence with a total complexity lower than that of the RSN [gower2019rsn].
@@
Both problems are widely used in practical machine learning applications. The logistic loss satisfies the strongly convex assumption under mild conditions [sypherd2020alpha], whereas the Poisson loss is a self-concordant function. Table 1 shows details of the datasets used and the associated loss function.
@@ 5.1 Implementation Details
Note that, under Definition 5.1, we have $\omega=\xi=1$ and thus the convergence rate for strongly convex functions is simplified as in (53). From now onward, unless otherwise specified, when Algorithm 1 uses fixed operators, the former definition of the $S_{n_{i}}$ subsets will be used to construct $R_{i}$ in Definition 5.1. We call this version of Algorithm 1 AML–Newton (fixed) as it uses fixed subspaces. If Algorithm 1 performs steps using iteration-dependent $R_{i,k}$, then they are generated based on the latter definition of the $S_{n_{i}}$ subsets. We call this version of Algorithm 1 AML–Newton (random) since the subspaces are generated randomly at each iteration. Both versions of Algorithm 1 calculate the damping parameter $\alpha_{k}$ using the Armijo rule.
It is well known that multilevel (or subspace) methods are particularly effective for problems that have a low-rank structure, e.g., their Hessian is nearly low-rank, or the second-order information is concentrated in the leading eigenvalues [ho2019newton, tsipinakis2024multilevel]. In contrast, when the curvature is spread across many directions, the first phase of multilevel methods can be slow, which may prevent the algorithm from entering the quadratic phase. In this case, larger values of $\sigma$ should be selected in order allow for more Newton steps. Similarly, one could randomly permute the $m$ levels at each iteration, introducing an approximately $1/m$ probability of selecting the Newton step. If the problem exhibits a low-rank structure, we wish the coarse models to be employed across the first phase, and hence $\sigma$ should be selected small, nevertheless not too small to compromise the quadratic second stage of the algorithm.